OpenStudy (anonymous):

5/11 + 3 (show steps, please & thank you)

OpenStudy (anonymous):

5/11 + 3/1 => 5/11 + 33/11 => 38/11 => 3 5/11

OpenStudy (anonymous):

Thank you @GT

OpenStudy (ash2326):

We have \[\frac{5}{11}+3\] multiply and divide 3 by 11 \[\frac{5}{11}+3\times \frac{11}{11}\] we get \[\frac{5}{11}+\frac{33}{11}\] \[\frac{5+33}{11}\] finally \[\frac{38}{11}\]

OpenStudy (anonymous):

Oh, thank you for the steps! @ash2326 !